OpenStudy (anonymous):

What is f(-3) if f(x)=|2x-1|-5? I don't know how to do this and my homework is taking all night every night :( A. 2 B. -12 C. 0 D. -10

OpenStudy (john_es):

You have to plug in the value x=-3 in the expression of f(x). But remember that the absolute value of a number is the positive value of the number. I mean, \[f(-3)=|2\cdot(-3)-1|-5=|-7|-5=7-5=2\]

OpenStudy (anonymous):

well, thank you. im sorry that i always ask stupid questions on here but i always take all night to do homework every night and still get behind. my therapist says that it is probably because my brain's math level is that of a middle schooler even though im 16. i really wish i could do all these things myself -.-